a research firm reports that the average annual salary in Nairobi area is $ 50,542.Suppose annual salaries in Nairobi area are normally distributed with a standard deviation of & 4,246.A Nairobi worker is randomly selected.

1)what is the probability that the workers annual salary is mire than & 60,000?

2)what is the probability that the workers annual salary is less than & 45,000

3) what is the probability that the workes annual salary is between & 44,000 and & 52,000?



1
Expert's answer
2021-11-26T11:56:51-0500

1) Probability of more than 60,000

"P(X>60000)=P(\frac{x-\mu}{\sigma} >\frac{60000-50542}{4246})"


"P(X>60000)=P(Z>2.23)"


"P(X>60000)=0.0129"


2) Probability of less than 45,000

"P(X<45000)=P(\frac{x-\mu}{\sigma} <\frac{45000-50542}{4246})"


"P(X<45000)=P(Z<-1.31)"


"P(X<45000)=1-0.0951=0.9049"


3) Probability of between 44,000 & 52,000

"P(44000<X<52000)=P(\frac{44000-50542}{4246}<\frac{x-\mu}{\sigma} <\frac{52000-50542}{4246})"


"P(44000<X<52000)=P(-1.54<Z <0.34)"


"P(44000<X<52000)=0.6331-0.0618"


"P(44000<X<52000)=0.5713"